Circular No. 6992

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dl IN NGC 1084 J. Y. King, M. Modjaz, T. Shefler, E. Halderson, W. D. Li, R. R. Tref-

fers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dl was found on an unfiltered image taken on Aug. 20.5 UT (mag ∼ 16.0), and was confirmed on earlier images taken on Aug. 2.5 (mag ∼ 16.0) and 9.5 (mag ∼ 15.9). The new object is located at α = 2^{h}46^{m}01^{s}.47, δ = −7^{o}34′25′′.1 (equinox 2000.0), which is ∼ 21′′.6 east and 11′′.6 north of the nucleus of NGC 1084. A KAIT image taken on 1997 Dec. 26.2 shows no star at this position (limiting mag ∼ 18.0). SN 1996an and SN 1963P also appeared in NGC 1084.

MXB 1730–335 D. W. Fox and W. H. G. Lewin, Massachusetts Institute of Technol-

ogy, report on behalf of a collaboration: “Observations with the All-Sky Monitor and Proportional Counter Array (PCA) of the Rossi X-ray Timing Explorer indicate the start of a new outburst of this rapid burster (cf. IAUC 6813). A PCA observation on Aug. 19.2 UT found the source emitting extremely bright (∼ 13 000 PCA counts/s, or ∼ 1 Crab), 100–200-s-long, type-II x-ray bursts, with persistent emission of ∼ 1000 PCA counts/s. Strong, 2–3-Hz, quasiperiodic oscillations (QPOs) were visible in one of the bursts, and strong 0.05-, 2–4-, and 0.4-Hz QPOs were visible at various times in the persistent emission, with the 0.4-Hz QPOs beginning 100–300 s before the bursts. Details, plots, and reports of observations are posted

at http://space.mit.edu/~derekfox/RB/. High-time-resolution (< 10 s)

infrared observations are particularly sought after.”

COMET C/1998 M5 (LINEAR) Total visual magnitude estimates: July 25.96 UT, 12.1 (M. Meyer,

Frauenstein, Germany, 0.25-m reflector); 31.01, 12.0 (K. Sárneczky, Ágasvár, Hungary, 0.44-m reflector); Aug. 8.88, 11.1 (M. Lehký, Hradec Králové, Czech Republic, 25×100 binoculars; moonlight); 13.28, 11.4 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 17.84, 11.5 (V. S. Nevski, Vitebsk, Belarus, 0.30-m reflector); 20.02, 11.7 (J. Carvajal, Toledo, Spain, 0.32-m reflector); 20.89, 10.7 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector).

1998 August 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green